OpenStudy (anonymous):

If \( y^3 = 16 x^2\), determine \(\frac{dx}{dt}\) when \( x = 4\) and \(\frac{dy}{dt} = -3 \) . \(\frac{dx}{dt} = ?\)

OpenStudy (anonymous):

sorry abt the format

OpenStudy (anonymous):

i dont know how to make it in better format

OpenStudy (ash2326):

Find \[\frac{dx}{dt}\ when\ \frac{dy}{dt}=-3\]

OpenStudy (anonymous):

thanks

OpenStudy (beginnersmind):

Differentiate y^3 = 16 x^2 with respect to t and plug in dy/dt.

OpenStudy (ash2326):

@rukh did you understand ? Can you do it ?

OpenStudy (across):

Let\[y^3=16x^2.\]Then\[3y^2\cdot y'=32x\cdot x',\]\[3y^2\cdot (-3)=32(4)\cdot x',\]\[-9y^2=128x',\]\[x'=-\frac{9}{128}y^2.\]

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

@ ash2326 yes i can try...thank u

OpenStudy (ash2326):

You're welcome but thank @across :D